OpenStudy (anonymous):

A set of test scores is normally distributed with a mean of 145 and a standard deviation of 11. Seif received a score of 167. What was his percentile score?

OpenStudy (anonymous):

\[P(X<167)=P\left(\frac{X-145}{11}<\frac{167-145}{11}\right)=P(Z<2)=k\] The value of \(k\) is the \(k\)th percentile.